# Dynamical Formation of Apparent Horizons due to Boundary Effect in Vacuum Einstein Gravity

### v2

@@ 4.5. Signature for decay rates and scale-invariant norms
Perhaps the most challenging aspect of trapped surface formation results, historically, has been the attempt to find initial data that are, in an appropriate sense, large (this is by necessity, as is implied by the monumental work of [19]) but also small enough to allow for an existence result of a spacetime region that gives trapped surfaces the time they would require to form. The first such initial data set, in the absence of symmetries, was given by [17]. Later contributions include [22], [6] and [2]. Moreover, one would have to construct norms which preserve, at least approximately, the hierarchy present in the initial data upon evolution of the Einstein equations. The signature for decay rates, which was first introduced in [3], is the tool we will use in the present paper to build scale-invariant norms. These will be norms that, upon evolution of the initial data, remain bounded above by a uniform constant (with the exception of a few anomalous terms). For another application of this framework, see [4].
@@
For any horizontal tensor-field $\phi$ and Yang-Mills component $\mathcal{Y}$, we define the following norms:
@@ 4.6. Norms
Let $N\geq 3$ be a natural number. Let $\psi_{g}\in\begin{Bmatrix}\omega,\mbox{tr}\chi,\hat{\chi},\underline{\omega},\eta,\underline{\eta}\end{Bmatrix}$, $\Psi_{u}\in\begin{Bmatrix}\beta,\rho,\sigma,\underline{\beta}\end{Bmatrix}$ and $\Psi_{\underline{u}}\in\begin{Bmatrix}a^{\frac{1}{2}}\rho,a^{\frac{1}{2}}\sigma,\underline{\beta},\underline{\alpha}\end{Bmatrix}$. Moreover, we will sometimes use $\Psi$ to denote an arbitrary $\Psi_{u}$ or a $\Psi_{\underline{u}}$. Also, define $\widetilde{\mbox{tr}\underline{\chi}}:=\mbox{tr}\underline{\chi}+\frac{2}{\lvert u\rvert}$. For $0\leq i\leq N$, we define
@@
(94) $+\frac{a}{\lvert u\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}+\frac{a}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\widetilde{\mbox{tr}\underline{\chi}}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})},$
$\mathcal{R}_{i,\infty}(u,\underline{u}):=\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{\underline{u}}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})},$
@@
(95) $\begin{split}\Gamma_{j,2}(u,\underline{u}):=&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\psi_{g}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+\frac{a^{\frac{1}{2}}}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\hat{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\\ &+\frac{a}{\lvert u\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+\frac{a}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\widetilde{\mbox{tr}\underline{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})},\end{split}$
(96) $\mathcal{R}_{i,2}(u,\underline{u}):=\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{\underline{u}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})},$
@@
(97) $\mathcal{R}_{i}(u,\underline{u}):=\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{u}}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})},$
(98) $\underline{\mathcal{R}}_{i}(u,\underline{u}):=\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta^{R}}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{\underline{u}}}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}$
@@ 5.1. Preliminary bootstrap assumptions
We will be employing a bootstrap argument to obtain a priori bounds on $\Gamma$ and $\mathcal{R}$. Along the initial hypersurfaces $H_{u_{\infty}}$ and $\underline{H}_{0}$, an analysis of the initial data using transport equations (see, for example, [6]) yields
$\Gamma_{0}+\mathcal{R}_{0}\lesssim\mathcal{I}.$
Our goal is to show that in the entire region
$\mathcal{D}:=\begin{Bmatrix}(u,\underline{u},\theta^{1},\theta^{2})\hskip 2.84526pt\mid\hskip 2.84526ptu_{\infty}\leq u\leq-a,\hskip 1.42262pt0\leq\underline{u}\leq 1\end{Bmatrix}$
there exists a constant $c(\mathcal{I})=\mathcal{I}^{4}+\mathcal{I}^{2}+\mathcal{I}+1$ such that
$\Gamma+\mathcal{R}\lesssim c(\mathcal{I}).$
We assume, as a bootstrap assumption, the following:
(102) $\Gamma\leq\Gamma,\hskip 5.69054pt\mathcal{R}\leq R,$
where $\Gamma$ and $\hskip 1.42262ptR$ are large so that
$\mathcal{I}^{4}+\mathcal{I}^{2}+\mathcal{I}+1\ll\text{min}\begin{Bmatrix}\Gamma,\hskip 1.42262ptR,\hskip 1.42262ptM\end{Bmatrix}$
but also such that
$(\Gamma+R)^{20}\leq a^{\frac{1}{16}}.$
5.2. Estimates on the metric components
For the metric component $\Omega$, the induced metric $\gamma$ of $S_{u,\underline{u}}$ and for the area of $S_{u,\underline{u}}$ the following propositions hold:
@@ 5.2. Estimates on the metric components
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), we have
$\lVert\Omega-1\rVert_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{\Gamma}{\lvert u\rvert}$.
@@
hold throughout the slab of existence $\mathcal{D}$.
@@
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), fix a point $(u,\theta)$ on the initial hypersurface $\underline{H}_{0}$. Let $\Lambda(u)$ and $\lambda(u)$ be the largest and smallest eigenvalues of $\gamma^{-1}(u,0,\theta)\hskip 1.42262pt\gamma(u,\underline{u},\theta)$ respectively, along the outgoing null geodesics emanating from $(u,\theta)$. There holds
@@
Proposition 5.4.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), for the area of the 22-sphere $S_{u,\underline{u}}$, there holds
$\sup_{\underline{u}}\lvert\text{Area}(S_{u,\underline{u}})-\text{Area}(S_{u,0})\rvert\lesssim\frac{\Gamma^{\frac{1}{2}}}{a^{\frac{1}{2}}}\lvert u\rvert^{2}.$
The proofs of Propositions 5.1 to 5.4 are the same as in [1]
@@ 5.3. Estimates for transport equations
(103) $\lVert{\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\lVert\mathcal{G}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime\prime}})}+\int_{\underline{u}^{\prime\prime}}^{\underline{u}}\lVert{\nabla_{3}\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
(104) $\lVert{\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\lVert\mathcal{G}\rVert_{L^{2}_{(sc)}(S_{u^{\prime\prime},\underline{u}})}+\int_{u^{\prime\prime}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{\nabla_{4}\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
@@
where we have used the fact that $\mathrm{e}^{\Gamma\lVert u^{-2}\rVert_{L_{u}^{1}}}\lesssim\mathrm{e}^{\epsilon\Gamma/a}\lesssim 1$.
@@ 5.4. Sobolev embedding
Proposition 5.7.
Under the assumption on the initial data and the bootstrap assumption (2.10), the isoperimetric constant obeys the following bound
(110) $I(S_{u,\underline{u}})\leq\frac{1}{\pi}$
for $u\in[u_{\infty},-a]$ and $\underline{u}\in[0,\epsilon]$.
@@
Fix a $u$. For $U_{\underline{u}}$ a subset of $S_{u,\underline{u}}$, denote by $U_{0}\subset S_{u,0}$ the backward image of $U_{\underline{u}}$ under the diffeomorphism generated by the equivariant vector field $L.$ Using Propositions 5.2 and 5.3 and their proof, we can obtain the estimates
$\frac{\text{Perimeter}(\partial U_{\underline{u}})}{\text{Perimeter}(\partial U_{0})}\geq\sqrt{\inf_{S_{u,0}}\lambda(\underline{u})}$
@@
$\frac{\text{Area}(U_{\underline{u}})}{\text{Area}(U_{0})}\leq\sup_{S_{u,0}}\frac{\det(\gamma_{\underline{u}})}{\det(\gamma_{0})},\hskip 5.69054pt\frac{\text{Area}(U_{\underline{u}}^{c})}{\text{Area}(U_{0}^{c})}\leq\sup_{S_{u,0}}\frac{\det(\gamma_{\underline{u}})}{\det(\gamma_{0})}.$
The conclusion then follows from the fact that $I(S_{u,0})=\frac{1}{2\pi}$ and the bounds from Propositions 5.2, 5.3 . ∎
@@
Proposition 5.8.
Let $(S,\gamma)$ be a Riemannian 22-manifold with the isoperimetric constant $I(S)$. Then the following Sobolev embedding holds for any $\mathcal{G}\in\Gamma(^{N}\otimes T^{*}S)$
[…]
for any $p\in(2,\infty)$.

### v4

@@ 3.5. Signature for decay rates and scale-invariant norms
Perhaps the most challenging aspect of trapped surface formation results, historically, has been the attempt to find initial data that are, in an appropriate sense, large (this is by necessity, as is implied by the monumental work of [21]) but also small enough to allow for an existence result of a spacetime region that gives trapped surfaces the time they would require to form. The first such initial data set, in the absence of symmetries, was given by [19]. Later contributions include [25], [6] and [2]. Moreover, one would have to construct norms that preserve, at least approximately, the hierarchy present in the initial data upon evolution of the Einstein equations. The signature for decay rates, which was first introduced in [3], is the tool we will use in the present paper to build scale-invariant norms. These will be norms that, upon evolution of the initial data, remain bounded above by a uniform constant (with the exception of a few anomalous terms). In particular, the dispersive estimates for the Ricci coefficients and the Weyl curvature components are restricted by their transport equations, and the signature of decay rates allows one to systematically obtain such estimates. For another application of this framework, see [4].
@@
For any horizontal tensor-field $\phi$, we define the following norms:
@@ 3.6. Norms
[…]
@@
(108) $+\frac{a}{\lvert u\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}+\frac{a}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\widetilde{\mbox{tr}\underline{\chi}}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}$
(109) $+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\hat{\chi},\omega,\underline{\omega})}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}$
$\mathcal{R}_{i,\infty}(u,\underline{u}):=a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{u}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\alpha}}\rVert_{L^{\infty}_{(sc)}(S_{u,\underline{u}})},$
@@
(110) $\begin{split}\Gamma_{j,2}(u,\underline{u}):=&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\psi_{g}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+\frac{a^{\frac{1}{2}}}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\hat{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\\ &+\frac{a}{\lvert u\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+\frac{a}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\widetilde{\mbox{tr}\underline{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\hat{\chi},\omega,\underline{\omega})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})},\end{split}$
(111) $\mathcal{R}_{i,2}(u,\underline{u}):=a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\beta,\underline{\beta},\rho,\sigma)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})},$
@@
(112) $\mathcal{R}_{i}(u,\underline{u}):=a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{u}}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})},$
(113) $\underline{\mathcal{R}}_{i}(u,\underline{u}):=a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\Psi_{\underline{u}}}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}$
@@ 4.1. Preliminary bootstrap assumptions
We derive the a priori estimates for the geometric and curvature norms by a bootstrap argument on the double–null development region
$D:=\Big\{(u,\underline{u},\theta^{1},\theta^{2})\ \Big|\ u_{\infty}\leq u\leq-a,\ \ 0\leq\underline{u}\leq 1\Big\}.$
Throughout, all implicit constants are universal and independent of $a$, depending only on fixed structural constants of the equations and on the number of derivatives $N$.
We employ the scale–invariant norms $\Gamma$ (Ricci coefficients) and $\mathcal{R}$ (curvature components) introduced in Section 3.6. These are defined as sums of weighted $L^{\infty}_{sc}$ and $L^{2}_{(sc)}$ norms of angular derivatives up to order $N+4$ and $N+3$, respectively. In particular, $\Gamma$ controls all Ricci coefficients
$\psi\in\{\mbox{tr}\chi,\hat{\chi},\mbox{tr}\underline{\chi},\underline{\hat{\chi}},\eta,\underline{\eta},\omega,\underline{\omega}\}$
in scale–invariant norms, while $\mathcal{R}$ controls the Weyl curvature components
$\Psi\in\{\alpha,\beta,\rho,\sigma,\underline{\beta},\underline{\alpha}\}$
through the corresponding null energy fluxes and supremum norms.
Initial bounds. Along the initial null hypersurfaces $H_{u_{\infty}}$ and $\underline{H}_{0}$, the data are prescribed so that the full hierarchy of Ricci coefficients and curvature components satisfies the dispersive scaling assumptions of Theorem 2.2. A direct analysis of the null constraint and transport equations along the initial hypersurfaces (cf. [6] and references therein) yields the quantitative bound
(117) $\Gamma_{0}+\mathcal{R}_{0}\;\lesssim\;\mathcal{I},$
where $\Gamma_{0},\mathcal{R}_{0}$ denote the initial values of the norms on $H_{u_{\infty}}\cup\underline{H}_{0}$ and $\mathcal{I}$ is the size of the prescribed data.
Bootstrap assumptions. We assume, on the spacetime region $D$, the bootstrap bounds
(118) $\Gamma\leq\Gamma_{\ast},\qquad\mathcal{R}\leq R_{\ast},$
for fixed constants $\Gamma_{\ast},R_{\ast}\geq 1$ to be chosen. These constants are taken sufficiently large so that
(119) $\mathcal{I}^{4}+\mathcal{I}^{2}+\mathcal{I}+1\;\ll\;\min\{\Gamma_{\ast},R_{\ast},M_{\ast}\},$
where $M_{\ast}$ denotes the corresponding metric norm bound, and at the same time satisfy the compatibility condition
(120) $(\Gamma_{\ast}+R_{\ast})^{20}\leq a^{1/16}.$
The smallness encoded in (120) ensures that all error terms produced by nonlinear interactions of Ricci and curvature components remain perturbative after integration in $u$ and $\underline{u}$.
The aim is to prove that the estimates implied by (118) can in fact be improved to
$\Gamma+\mathcal{R}\;\lesssim\;c(\mathcal{I}),\qquad c(\mathcal{I}):=\mathcal{I}^{4}+\mathcal{I}^{2}+\mathcal{I}+1,$
throughout $D$. By a standard continuity argument, this yields a closed a priori bound and therefore semi–global control of the solution.
Estimates on the metric components
We first control the metric quantities $\Omega$, the induced sphere metric $\gamma$, and the area radius of $S_{u,\underline{u}}$. We begin with the null lapse $\Omega$.
@@
Under the assumptions of Theorem 2.2 and the bootstrap bounds (118), the null lapse satisfies, for every sphere $S_{u,\underline{u}}\subset D$,
$\|\Omega-1\|_{L^{\infty}(S_{u,\underline{u}})}\;\lesssim\;\frac{\Gamma_{\ast}\,a^{1/2}}{|u|^{2}}.$
Proof.
In double–null gauge the lapse obeys the transport equation
(121) $\nabla_{3}\log\Omega=-\,\underline{\omega}.$
On the initial outgoing hypersurface $H_{u_{\infty}}$ the data normalization gives
$\Omega(u_{\infty},\underline{u},\theta)=1\quad\text{for all }\underline{u}\in[0,1],$
hence $\log\Omega=0$ there. Integrating (121) along the incoming null generators from $u_{\infty}$ to $u$ at fixed $(\underline{u},\theta)$ yields
$\log\Omega(u,\underline{u},\theta)=-\int_{u_{\infty}}^{u}\underline{\omega}(u^{\prime},\underline{u},\theta)\,du^{\prime}.$
By the bootstrap bound on Ricci coefficients and the scale–invariant weights,
$|\underline{\omega}(u^{\prime},\underline{u},\theta)|\;\lesssim\;\frac{\Gamma_{\ast}\,a^{1/2}}{|u^{\prime}|^{3}}.$
Therefore,
$|\log\Omega(u,\underline{u},\theta)|\;\lesssim\;\Gamma_{\ast}a^{1/2}\int_{u_{\infty}}^{u}\frac{du^{\prime}}{|u^{\prime}|^{3}}\;\lesssim\;\frac{\Gamma_{\ast}a^{1/2}}{|u|^{2}}.$
Since the right–hand side is $\ll 1$ by (120), we conclude
$|\Omega(u,\underline{u},\theta)-1|\;\lesssim\;\frac{\Gamma_{\ast}a^{1/2}}{|u|^{2}}.$
Taking the supremum over $S_{u,\underline{u}}$ gives the stated estimate. ∎
@@
hold throughout the slab of existence $D$.
@@
Under the assumptions of Theorem 2.2 and the bootstrap assumptions (118), fix a point $(u,\theta)$ on the initial hypersurface $H_{\infty}$. Let $\Lambda(u)$ and $\lambda(u)$ be the largest and smallest eigenvalues of $\gamma^{-1}(u_{\infty},\underline{u},\theta)\hskip 1.42262pt\gamma(u,\underline{u},\theta)$ respectively, along the incoming null geodesics emanating from $(\underline{u},\theta)$. There holds
@@ 4.2. Estimates for transport equations
(122) $\lVert{\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\lVert\mathcal{G}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime\prime}})}+\int_{\underline{u}^{\prime\prime}}^{\underline{u}}\lVert{\nabla_{4}\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
(123) $\lVert{\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\lVert\mathcal{G}\rVert_{L^{2}_{(sc)}(S_{u^{\prime\prime},\underline{u}})}+\int_{u^{\prime\prime}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{\nabla_{3}\mathcal{G}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
@@
where we have used the fact that $\mathrm{e}^{\Gamma\lVert u^{-2}\rVert_{L_{u}^{1}}}\lesssim\mathrm{e}^{\Gamma/a}\lesssim 1$.
@@ 4.3. Sobolev embedding
Proposition 4.6 (Uniform control of the isoperimetric constant of $S_{u,\underline{u}}$).
Assume the initial data hypotheses and the bootstrap bounds (2.10) for the double–null development. Then for every sphere of the foliation
$S_{u,\underline{u}},\qquad u\in[u_{\infty},-a],\ \underline{u}\in[0,1],$
the isoperimetric constant satisfies the uniform estimate
(129) $I(S_{u,\underline{u}})\;\leq\;\frac{1}{\pi}.$
Here $I(S,\gamma)$ denotes the isoperimetric constant of the Riemannian 22–sphere $(S,\gamma)$, defined by
$I(S,\gamma):=\sup_{U\subset S}\frac{\min\{\text{Area}(U),\text{Area}(U^{c})\}}{\text{Per}(\partial U)^{2}},$
where the supremum ranges over all domains $U\subset S$ with smooth boundary.
@@
We first note that the isoperimetric constant is scale invariant: if $\gamma^{\prime}=\lambda^{2}\gamma$ on a surface $S$, then
$\frac{\min\{\text{Area}_{\gamma^{\prime}}(U),\text{Area}_{\gamma^{\prime}}(U^{c})\}}{\text{Per}_{\gamma^{\prime}}(\partial U)^{2}}=\frac{\lambda^{2}\min\{\text{Area}_{\gamma}(U),\text{Area}_{\gamma}(U^{c})\}}{\lambda^{2}\text{Per}_{\gamma}(\partial U)^{2}},$
hence $I(S,\gamma^{\prime})=I(S,\gamma)$. Therefore it suffices to prove the estimate for the renormalized metric
$\tilde{\gamma}_{u,\underline{u}}:=|u|^{-2}\gamma_{u,\underline{u}}\quad\text{on }S_{u,\underline{u}}.$
Suppose a metric $\gamma$ on $S^{2}$ satisfies, for some $\varepsilon\in(0,1)$,
(130) $(1-\varepsilon)\gamma_{0}\leq\gamma\leq(1+\varepsilon)\gamma_{0}$
as quadratic forms, where $\gamma_{0}$ is the unit round metric. Then for every smooth domain $U\subset S$,
$\text{Area}_{\gamma}(U)\leq(1+\varepsilon)\text{Area}_{\gamma_{0}}(U),\qquad\text{Area}_{\gamma}(U^{c})\leq(1+\varepsilon)\text{Area}_{\gamma_{0}}(U^{c}),$
@@
$\text{Per}_{\gamma}(\partial U)\geq(1-\varepsilon)^{1/2}\text{Per}_{\gamma_{0}}(\partial U),\quad\Rightarrow\quad\text{Per}_{\gamma}(\partial U)^{2}\geq(1-\varepsilon)\text{Per}_{\gamma_{0}}(\partial U)^{2}.$
Hence
$\frac{\min\{\text{Area}_{\gamma}(U),\text{Area}_{\gamma}(U^{c})\}}{\text{Per}_{\gamma}(\partial U)^{2}}\leq\frac{1+\varepsilon}{1-\varepsilon}\frac{\min\{\text{Area}_{\gamma_{0}}(U),\text{Area}_{\gamma_{0}}(U^{c})\}}{\text{Per}_{\gamma_{0}}(\partial U)^{2}}.$
Taking the supremum gives
(131) $I(S,\gamma)\leq\frac{1+\varepsilon}{1-\varepsilon}I(S,\gamma_{0}).$
For the unit round sphere one has $I(S^{2},\gamma_{0})=\frac{1}{2\pi}$. Thus if $\varepsilon\leq\frac{1}{3}$, then
(132) $I(S,\gamma)\leq\frac{1+\varepsilon}{1-\varepsilon}\frac{1}{2\pi}\leq\frac{1}{\pi}.$
In the double–null foliation the induced metric satisfies the transport equations
$\nabla_{4}\gamma_{AB}=2\chi_{AB},\qquad\nabla_{3}\gamma_{AB}=2\underline{\chi}_{AB},$
that is,
$\nabla_{4}\gamma_{AB}=(\mbox{tr}\chi)\gamma_{AB}+2\hat{\chi}_{AB},\qquad\nabla_{3}\gamma_{AB}=(\mbox{tr}\underline{\chi})\gamma_{AB}+2\underline{\hat{\chi}}_{AB}.$
Define $\tilde{\gamma}_{AB}=|u|^{-2}\gamma_{AB}$. Using $\nabla_{4}u=0$ and the standard normalization of $\nabla_{3}u$, one obtains schematically
(133) $\nabla_{4}\tilde{\gamma}_{AB}$ $=\Big(\mbox{tr}\chi-\frac{2}{|u|}\Big)\tilde{\gamma}_{AB}+2|u|^{-2}\hat{\chi}_{AB},$
(134) $\nabla_{3}\tilde{\gamma}_{AB}$ $=\Big(\mbox{tr}\underline{\chi}+\frac{2}{|u|}\Big)\tilde{\gamma}_{AB}+2|u|^{-2}\underline{\hat{\chi}}_{AB}+\text{l.o.t.}$
By the initial data assumptions on $H_{u_{\infty}}$ we have
$\tilde{\gamma}(u_{\infty},0)=\gamma_{0}+O(a^{-1/2}|u_{\infty}|^{-3}).$
The bootstrap bounds (2.10) give, uniformly for $u\in[u_{\infty},-a]$, $\underline{u}\in[0,1]$,
$\Big|\mbox{tr}\chi-\frac{2}{|u|}\Big|\lesssim a^{-1/2}|u|^{-1},\qquad|\hat{\chi}|\lesssim a^{-1/2}|u|^{-1},$
$\Big|\mbox{tr}\underline{\chi}+\frac{2}{|u|}\Big|\lesssim|u|^{-2},\qquad|\underline{\hat{\chi}}|\lesssim a^{1/2}|u|^{-2}.$
Since $\underline{u}\in[0,1]$ and $|u|\geq a\gg 1$, Grönwall’s inequality applied to (133) along the $\nabla_{4}$ direction and to (134) along the $\nabla_{3}$ direction yields
(135) $\sup_{u\in[u_{\infty},-a],\ \underline{u}\in[0,1]}\|\tilde{\gamma}_{u,\underline{u}}-\gamma_{0}\|_{C^{0}(S_{u,\underline{u}})}\lesssim a^{-1/2}.$
For $a\geq a_{0}$ sufficiently large, this implies the bilipschitz comparison (130) with some $\varepsilon\leq\frac{1}{3}$, uniformly for all $S_{u,\underline{u}}$ in the slab.
By scale invariance,
$I(S_{u,\underline{u}},\gamma_{u,\underline{u}})=I(S_{u,\underline{u}},\tilde{\gamma}_{u,\underline{u}}).$
Combining (135) with (132) gives
$I(S_{u,\underline{u}})\leq\frac{1}{\pi}$
uniformly for all $u\in[u_{\infty},-a]$, $\underline{u}\in[0,1]$. This proves the proposition. ∎
@@
Proposition 4.7 (Sobolev inequality for tensorfields in terms of the isoperimetric constant).
Let $(S,\gamma)$ be a smooth compact Riemannian 22–manifold. Denote by $\text{Area}(S)$ its $\gamma$–area and by $I(S)$ its isoperimetric constant
$I(S):=\sup_{U\subset S}\frac{\min\{\text{Area}(U),\text{Area}(U^{c})\}}{\text{Per}(\partial U)^{2}},$
where the supremum is taken over all domains $U\subset S$ with smooth boundary. Then for every $p\in(2,\infty)$ there exists a constant $C_{p}<\infty$ depending only on $p$ such that the following holds.
Let $\mathcal{G}$ be any smooth tensorfield of type $(0,N)$ on $S$ (in particular, $\mathcal{G}\in\Gamma(\otimes^{N}T^{*}S)$; more generally the same statement holds for any tensor bundle equipped with the metric induced by $\gamma$). Then
(136) $\text{Area}(S)^{-\frac{1}{p}}\,\|\mathcal{G}\|_{L^{p}(S)}\;\leq\;C_{p}\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\Big(\|\nabla\mathcal{G}\|_{L^{2}(S)}+\text{Area}(S)^{-\frac{1}{2}}\,\|\mathcal{G}\|_{L^{2}(S)}\Big).$